2. Solve the initial value problem = 4x + 2y, = 2x + y with r(0) = 1, y(0) = 0.
Answer #1

Let A = { 1 2 Put 1 Anal=0>14 21 12 il (4-x)(1-x)-450 > ²_5x=0 5x=0,5 are eigenvolves. for 7,20 u 27 sono Let d=1&p=-2 7 is eه[3] اعلام و الم ) [ Hena | +[4] >= 3*د 0 = (0)۱ , ۶ (۵) ۷ ارم = ]H5[ 32 [1] = دو ر =م) * 0 from Answer کمار ۲۶ کم و م و =